A hiummingbird can flutter it’s wings 4,800 times per minute what is the frequency of wing flutters per second? what is the period for one wing in seconds? Round answer to nearest decimal
1
Expert's answer
2018-06-13T08:22:08-0400
a) We can find the frequency of wing flutters per second as follows:
f=(4800 times)/(1 min)∙(1 min)/(60 s)=80 Hz.
b) Finally, we can find the period for one wing:
T=1/f=1/(80 Hz)=0.0125 s.